The greater of two numbers is 1 more than twice the smaller. Three times the greater exceeds 5 times the smaller by 10. Find the numbers.
Answer by oberobic(2304)

a = greater of 2 numbers
b = lesser of 2 numbers
.
a = 2b + 1
.
3*a = 5*b + 10
.
substitute for a
.
3(2b+1) = 5b +10
6b + 3 = 5b + 10
b = 7
.
a = 2(7)+1
a = 15
.
Done.